### v1

@@ 5.3 Statistical analysis
Some statistical tests are conducted to further analyze the computational results we received in Section 5.2. We first conduct the Friedman test demvsar2006statistical to evaluate the classification accuracy. The null hypothesis is that all the tested models have the same classification accuracy.
Let $p$ and $q$ be the number of datasets and the number of tested models, respectively. Hence, $p=20$ and $q=9$, and the $\chi_{F}^{2}$ value of the Friedman test can be calculated by the formula (4)
$\chi_{F}^{2}=\frac{12p}{q(q+1)}\left(\sum_{i=1}^{q}R_{i}^{2}-\frac{q(q+1)^{2}}{4}\right)\approx 73.78,$ (4)
@@
$F_{F}=\frac{(p-1)\chi_{F}^{2}}{p(q-1)-\chi_{F}^{2}}\approx 15.66,$ (5)
where $F_{F}$ is distributed according to the $F$-distribution with degrees of freedom $(8,160)$. With the significance level to be $\alpha=0.05$, the critical value of $F_{F}$ in our case is about 2.03, which is far less than the value of $F_{F}$. Therefore, we are confident in rejecting the null hypothesis. In other words, there exists a significant difference among all the tested models.
@@
$CD=q_{.05}\sqrt{\frac{q(q+1)}{6p}}\approx 2.62,$ (6)
where $q_{.05}=3.10$ is the critical value of Tukey distribution. Therefore, the Nemenyi post-hoc test detects a significant difference between the proposed Im-LS-$\mathfrak{U}$-QTSVM model and each other tested model except the LS-$\mathfrak{U}$-QTSVM. The result from the Nemenyi post-hoc test is visualized in Figure 3.
@@
Even though the average ranks of the proposed model and LS-$\mathfrak{U}$-QTSVM are close, we may still find the difference by the Wilcoxon signed-rank test as suggested in benavoli2016should . The null hypothesis is that there is no significant difference between the accuracy of the two models. Let $\beta_{j}$ be the difference between the accuracy of the two models on the $j$th dataset. The calculated values $R^{+}$ and $R^{-}$ defined as follows
$R^{+}$ $=\sum_{\beta_{i}>0}\text{rank}(\beta_{i})+\frac{1}{2}\sum_{\beta_{i}=0}\text{rank}(\beta_{i})=185,$
$R^{-}$ $=\sum_{\beta_{i}<0}\text{rank}(\beta_{i})+\frac{1}{2}\sum_{\beta_{i}=0}\text{rank}(\beta_{i})=25,$
Let $R_{\min}=\min\{R^{+},R^{-}\}=25$ and calculate the statistics $z$:
$z=\frac{R_{\min}-\frac{1}{4}p(p+1)}{\sqrt{\frac{1}{24}p(p+1)(2p+1)}}\approx-3.15<-1.96.$
Therefore, the null hypothesis can be rejected with a $0.05$ confidence level.
@@
6 Conclusions
In this paper, we introduced three kernel-free Universum quadratic surface support vector machine models for binary classification: $\mathfrak{U}$-QTSVM, LS-$\mathfrak{U}$-QTSVM, and Im-LS-$\mathfrak{U}$-QTSVM. Binary classification with imbalanced classes presents significant challenges in machine learning, especially when the minority class is underrepresented. Traditional classifiers often fail to accurately capture the characteristics of the minority class, leading to biased models with poor predictive performance. To address this issue, we developed the Im-LS-$\mathfrak{U}$-QTSVM model, tailored for imbalanced datasets. We examined several theoretical properties and proposed an efficient algorithm for this model. Our computational experiments confirmed the effectiveness and efficiency of the proposed models. The main findings are summarized below:
  • We incorporated Universum data into kernel-free QTSVM models to propose the $\mathfrak{U}$-QTSVM model, followed by its least-square version, LS-$\mathfrak{U}$-QTSVM.
  • To tackle the challenge of imbalanced datasets, we adapted $\mathfrak{U}$-QTSVM to create Im-$\mathfrak{U}$-QTSVM and its least-square version, Im-LS-$\mathfrak{U}$-QTSVM, which reduces computational complexity by solving a system of equations.
@@ 6 Conclusions
Future research could extend this work in several ways. Immediate future work includes investigating the robustness of the proposed models for binary classification. Additionally, these models could be extended to semi-supervised learning scenarios and generalized for multi-class classification, which warrants further investigation.
@@
  • Informed consents: Informed consent was obtained from all individual participants included in the study.

### v2

@@ 5.3 Statistical analysis
Some statistical tests are conducted to further analyze the computational results we received in Section 5.2. We first conduct the Friedman test demvsar2006statistical to evaluate the classification accuracy in Table 3. The null hypothesis is that all the tested models have the same classification accuracy.
Let $p$ and $q$ be the number of datasets and the number of tested models, respectively. Hence, $p=21$ and $q=12$, and the $\chi_{F}^{2}$ value of the Friedman test can be calculated by the formula (4)
$\chi_{F}^{2}=\frac{12p}{q(q+1)}\left(\sum_{i=1}^{q}R_{i}^{2}-\frac{q(q+1)^{2}}{4}\right),$ (4)
@@
$F_{F}=\frac{(p-1)\chi_{F}^{2}}{p(q-1)-\chi_{F}^{2}},$ (5)
where $F_{F}$ is distributed according to the $F$-distribution with degrees of freedom $(q-1,(p-1)\times(q-1))=(11,220)$. Similarly, the Friedman test is applied to evaluate the G-mean recorded in Table 4 using formulas (4) and (5). The key parameters of Friedman tests on accuracy and G-mean are listed in Table 6. With the significance level to be $\alpha=0.05$, the critical value of $F_{F}$ is about 2.03, which is far less than the $F_{F}$ of either accuracy or G-mean. Therefore, we are confident in rejecting the null hypothesis. In other words, there exists a significant difference among all the tested models with respect to classification accuracy and G-mean.
Metrics $\chi_{F}^{2}$ $F_{F}$
Accuracy 109.385 17.990
G-mean 106.308 17.051
Table 6: Key parameters of Friedman test.
@@
$CD=q_{.05}\sqrt{\frac{q(q+1)}{6p}}\approx 3.636,$ (6)
where $q_{.05}=3.268$ is the critical value of the Tukey distribution. The results from the Nemenyi post-hoc test on both accuracy and G-mean are visualized in Figure 3.
(a) Nemenyi test on Accuracy.
(b) Nemenyi test on G-mean.
@@
From Figure 3(a), we notice that the average rank in accuracy of Im-LS-$\mathfrak{U}$-QTSVM is close to LS-$\mathfrak{U}$-QTSVM. From Figure 3(a), we notice that the average rank in G-mean of Im-LS-$\mathfrak{U}$-QTSVM is close to CSSVM. Therefore, we further apply the Wilcoxon signed-rank test as suggested in benavoli2016should to show the difference between the Im-LS-$\mathfrak{U}$-QTSVM model and each other tested model. For each metric, the null hypothesis is that there is no significant difference between the two models. Let $\beta_{j}$ be the difference between the metric of the two models on the $j$th dataset. The calculated values $R^{+}$ and $R^{-}$ defined as follows
$R^{+}$ $=\sum_{\beta_{i}>0}\text{rank}(\beta_{i})+\frac{1}{2}\sum_{\beta_{i}=0}\text{rank}(\beta_{i}),$
$R^{-}$ $=\sum_{\beta_{i}<0}\text{rank}(\beta_{i})+\frac{1}{2}\sum_{\beta_{i}=0}\text{rank}(\beta_{i}),$
Let $R_{\min}=\min\{R^{+},R^{-}\}$ and calculate the statistics $z$:
$z=\frac{R_{\min}-\frac{1}{4}p(p+1)}{\sqrt{\frac{1}{24}p(p+1)(2p+1)}}.$
For each metric, we calculate the $z$ value between the Im-LS-$\mathfrak{U}$-QTSVM model and each benchmark model. All the $z$ values are recorded in Table 7. Notice that all the $z$-statistic values are less than 1.96-1.96, which leads to the rejection of the null hypothesis with a $0.05$ confidence level.
@@
Compared Models Accuracy G-mean
LS-$\mathfrak{U}$-QTSVM -3.15 -3.74
LS-$\mathfrak{U}$-TSVM -3.53 -2.94
LS-QTSVM -4.01 -4.01
LS-TSVM -4.01 -4.01
$\mathfrak{U}$-QTSVM -4.01 -4.01
$\mathfrak{U}$-TSVM -2.80 -3.89
TSVM -4.01 -4.01
SVM -3.91 -4.01
CSSVM -4.01 -2.94
URKWELM -3.32 -3.94
FHTPSVM -3.82 -3.94
Table 7: (Wilcoxon signed-rank test) $z$-statistic between Im-LS-$\mathfrak{U}$-QTSVM and each compared model.
6 Conclusions and Future Research Directions
In this paper, we introduced three kernel-free Universum quadratic surface support vector machine models for binary classification: $\mathfrak{U}$-QTSVM, LS-$\mathfrak{U}$-QTSVM, and Im-LS-$\mathfrak{U}$-QTSVM. Binary classification with imbalanced classes remains a major challenge in machine learning, especially when minority class instances are scarce. Traditional classifiers often exhibit bias toward the majority class, leading to poor prediction performance on the minority class and potentially serious consequences in real-world applications such as medical diagnosis, fraud detection, or rare-event prediction. To address this issue, we developed the Im-LS-$\mathfrak{U}$-QTSVM model, tailored for imbalanced datasets. We examined several theoretical properties and proposed an efficient algorithm for this model. Our computational experiments confirmed the effectiveness and efficiency of the proposed models. The main findings are summarized below:
  • We proposed the $\mathfrak{U}$-QTSVM model, which leverages Universum points to provide additional prior knowledge that supports minority class representation. This approach enhances the classifier’s ability to capture subtle patterns in complex datasets, improving generalization performance without requiring kernel-based transformations. Its least-squares version, LS-$\mathfrak{U}$-QTSVM, further simplifies the optimization problem by converting it into a system of linear equations, making it computationally efficient for moderate-scale datasets.
  • To tackle the challenge of imbalanced datasets, we adapted $\mathfrak{U}$-QTSVM to create Im-$\mathfrak{U}$-QTSVM and its least-square version, Im-LS-$\mathfrak{U}$-QTSVM, which reduces computational complexity by solving a system of equations. Indeed, these models explicitly focus on improving classification of the minority class by adjusting the margin and leveraging Universum points strategically, which reduces the bias toward the majority class. The least-squares formulation reduces computational complexity while maintaining strong predictive performance, making it suitable for real-world imbalanced datasets.
@@ 6 Conclusions and Future Research Directions
Building on this work, future research could extend the proposed models to multi-class imbalanced datasets to handle more complex classification scenarios. Exploring adaptive selection or generation of Universum points may further improve generalization and reduce the need for manual parameter tuning. Developing efficient solvers or approximation methods for high-dimensional or large-scale datasets would enhance the models’ practical applicability. Immediate future work also includes assessing the robustness of the proposed models across diverse binary classification tasks and exploring extensions to semi-supervised learning settings.
In summary, the proposed Universum quadratic surface SVM models provide a flexible, computationally efficient, and effective solution for imbalanced binary classification. By addressing the limitations of traditional SVMs and explicitly considering minority class support, these models open up opportunities for reliable deployment in real-world applications. The future directions outlined above provide a roadmap for enhancing the generalizability, scalability, and robustness of these models, making them even more compelling for researchers and practitioners.
@@
  • Informed consent: Informed consent was obtained from all individual participants included in the study.
  • Data availability statement: The following GitHub repository provides all the links to the publicly available datasets that have been used for the numerical experiments of this study: https://github.com/tonygaobasketball/Sparse-UQSSVM-Mode ls-for-Binary-Classification. The code will be provided by the authors upon request.